40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6024     Entry time: Mon Nov 28 15:00:20 2011
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y arm ALS engaged 
Quote from #5894

  However I couldn't close the ALS loop somehow.

 Locking activity last night:

  It became able to close the ALS loop (beat-note signal was fed back to ETMY).

The UGF was about 60 Hz, but somehow I couldn't bring the UGF higher than that.

Every time when I increased the UGF more than 60 Hz, the Y end PDH was unlocked (or maybe ETMY became crazy at first).

Perhaps it could be a too much noise injection above 60 Hz, since I was using the coarse frequency discriminator.

Anyway I will try a cavity sweep and the successive noise budgeting while holding the arm length by the beat-note signal.

Another thing : I need a temperature feedback in the Y end green PDH loop, so that the PZT voltage will be offloaded to the laser temperature.

ELOG V3.1.3-